# Low Temperature Properties of Low-Loss Macroscopic Lithium Niobate Bulk Acoustic Wave Resonators

### v2

@@ I Introduction
Figure 1: a) Diagram of experimental set up. Cryostat environment was held under vacuum at temperatures of either 300 K or 4 K, and 10mK depending on the required measurement. b) Simulated mode distribution of the fifth order longitudinal mode. Red (blue) denotes the positive (negative) displacement in the $z$ direction
Piezoelectric lithium niobate for example provides the vibrating media for tunable micro-ring resonators [22, 23], optical frequency combs [24], micro electromechanical system (MEMS) based resonators [25], devices for optical to microwave transduction [26, 27, 28, 29, 30] and acoustic quantum state operations [31, 32, 33]. This material presents an attractive candidate as a macroscopic BAW resonator due to a stronger piezoelectric coupling coefficient, and smaller loss tangent than crystalline quartz at room temperature. However, material properties display a strong temperature dependence, and are thus drastically different when cooled. This motivates further investigation into the performance of macroscopic lithium niobate BAWs at cryogenic temperatures, for the purpose of next generation tests of fundamental physics. Additionally, the study of BAW modes at cryogenic temperatures allows one to investigate fundamental properties of the material.
In this work the loss mechanisms of lithium niobate BAW resonators are investigating by precise measurement of the quality factor ($Q$) of their acoustic modes at cryogenic temperatures. In the following sections II and III, the measurement set up is introduced and the frequencies of various fundamental bulk vibration modes are identified. Section IV then derives a theoretical model for the total $Q$ of a BAW resonator subject to accumulative loss mechanisms. In sections V and VI the measured $Q$ of various resonator modes are presented as functions of frequency and temperature respectively. Multiple competing loss mechanisms are identified based upon these observations. In section VII, we report on a strong drive amplitude dependence on the measured losses. Such self induced absorption and transparency effects are confirmed to be due to microscopic impurities by the use of an appropriate model of non-linear dissipation. Lastly, section VIII discusses the performance of these devices when further cooled to the extreme sub-Kelvin temperature of 10 mK.
@@ II Measurement Set Up
In this configuration, as shown in figure 1, acoustic resonances can be excited by an electrical drive signal applied to the electrodes. By connecting each co-axial line to ports of a vector network analyser; the acoustic mode frequencies $f_{r}$ are identified as peaks of maximum power transmission $S_{21}$. The input drive signal was varied across the frequency spectrum with 1 Hz resolution such that narrow resonances can be identified. The resulting spectra are then compared to a finite element model such that the mode structure of strongly coupled resonances can be identified.
Figure 2: Measurement of a)$f\cdot Q$ product and b) intrinsic $Q$ as a function of frequency for various longitudinal modes in the crystal resonators at $T=4$ K. The measured intrinsic quality factors are represented in blue, the black trace and grey bands denote the best fit and 95% confidence interval to the model of equation (3), the red trace shows the idealistic performance attainable without the presence of Rayleigh scattering. And the dashed black line gives the intrinsic limit to losses observed in similar quartz devices [35].
@@ III Mode Structure
The frequencies of strongly coupled modes found by measurement were matched to the frequencies given by solutions to a finite element modelling as in figure 1 b), the results of which are presented in table 1. Close agreement between experimental data and model solutions is found. All odd overtones from $n=3$, to $n=15$, in both A and B polarisation are recorded in the table. For higher frequencies, the finite element resolution of the model needs to be further increased in order to guarantee accuracy. However from the results provided in table 1, a simple linear extrapolation $f_{i}/n_{i}=f_{i}/n_{i}$ can be utilised to identify any $n_{ij}$ overtone, of frequency $f_{ij}$, for modes of the same polarisation family.
@@ IV Theoretical Model of Acoustic Loss
Figure 3: $Q$ as a function of temperature $T$ is measured for the B13,0,0 mode and the resulting scaling relationship between losses $1/Q$ and temperature is shown. Inset shows the scaling for the A19,0,0 mode when cooled to 10 mK. Temperature scaling characteristic of TLS limited loss is identified, however the presence of strong absorption peaks in close proximity are also observed.
The dominant loss mechanisms in BAW resonators can be understood by measurement of the intrinsic quality factor of the acoustic modes $Q_{i}$. This quantity is the inverse of the total phonon losses and as such can be parameterised as the inverse sum of the quality factor associated with each competing loss mechanism. For bulk waves, the contributing losses to $Q_{i}$ can be due to clamping of the vibrational structure to external supports, thermoelastic damping, phonon-phonon scattering, and phonon-defect scattering. Phonon-electron scattering loss can be dominant in semi-conductor systems or at microwave frequencies, however it is negligible here.
For acoustic resonators at 300 K, phonon-phonon scattering loss is well described by the macroscopic Akheiser regime [36], which predicts a constant $Q\cdot f$ product. For lithium niobate the theoretical limit to $Q\cdot f$ as predicted by the Akheiser model is $Q\cdot f\approx 4.74\times 10^{14}$, although the largest room temperature value observed in the devices measured in this work is $1.12\times 10^{14}$ corresponding to a maximum $Q$ of $3.28\times 10^{6}$ However, at lower temperatures the life time of thermal phonons $\tau_{\mathrm{th}}$ increases such that phonon attenuation process undergoes a phase change into the microscopic description of Landau and Rumer [37]. In this regime, the attenuation displays a weaker frequency dependence. As a result $Q\cdot f$ begins to scale as $Q\cdot f\propto f$, and $Q$ also greatly increases with decreasing temperature. It is precisely this change to the fundamental limit of intrinsic acoustic loss that gives rise to the numerous applications of high overtones BAW devices at low temperature. For lithium niobate at 4 K phonon modes are well described by the Landau-Rumer model for $f\gtrsim 1.4$ MHz, which equivalently gives $12\pi f\tau_{\mathrm{th}}>$ 1.
@@
where $\rho$ is the material density, $T$ is the reference temperature, $C^{s}$ is the specific heat capacity at constant strain, $c_{ij}$ is the isothermal elastic tensor and $\alpha_{ij}$ the thermal expansion tensor. The turning point frequency $f^{*}$ denotes the transition point between the adiabatic and isothermal propagation regimes, it is given by $f^{*}=V_{3}^{2}/2\pi\chi_{3}$ where $\chi_{3}$ is the material diffusivity. For lithium niobate at $T=4$ K the thermal conductivity and specific heat have been measured, but the elastic constants and expansion coefficients are unknown. Approximating the elastic constants from their room temperature values gives $f^{*}\approx 13$ MHz, with $f>f^{*}$ defining the isothermal regime.
@@
Acoustic attenuation due to either structural defects or microscopic impurities can be harder to quantify exactly, however scaling relationships can be predicted in some cases. Considering the standard tunnelling model of two level systems in amorphous glasses [40, 41], it can be shown [42] that the acoustic absorption by two level systems, potentially arising from non-crystalline impurities, results in a temperature dependent scaling of the acoustic losses $1/Q\propto T^{1/3}$. At lower temperatures populations of these two level systems can become unsaturated and have a dominant impact on total acoustic dissipation.
In addition to temperature scaling, Rayleigh-type scattering loss of phonons in metals and glasses is predicted to scale as $1/Q_{\mathrm{Rayleigh}}\propto f^{-3}$ [43]. This relationship however, has also been observed for high overtones in similar quartz piezoelectric BAW systems [35], suggesting the presence of small populations of impurity ions that cause amorphous behaviour at high frequencies.
@@
$\frac{1}{Q_{i}}=\frac{1}{Q_{\mathrm{Clamping}}}+\frac{1}{Q_{\mathrm{TED}}}+\frac{1}{Q_{\mathrm{Rayleigh}}}.$ (3)
@@ V Frequency Dependence of Acoustic Loss
The intrinsic quality factor of bulk lithium niobate modes can be measured by fitting a Lorentzian model to the $S_{21}$ spectrum of a given mode and extracting its linewidth $\gamma$, such that $Q_{i}=f_{r}/\gamma$. The measurement uncertainty following this procedure is then quantified as the standard error associated to the fit parameters. As the motional resistance of these modes are far greater than the source impedance of the electrical measurement network, the modes are severely under coupled, and thus the measured $Q$ corresponds exactly to the intrinsic quality factor of the resonator $Q_{i}$. The measured values from this procedure are shown in figure 2. The cryogenic quality factors show consistent improvement from room temperature values across all frequencies, with a maximum recorded quality factor of $8.9\times 10^{6}$.
By substituting equations (1) and (2) and the functional form $Q_{\mathrm{Rayleigh}}=10^{a}f^{-3}$ into equation (3), a fit can be performed to the measured quality factors as a function of frequency with $\lambda_{33}$, $\eta$ and $a$ as fit parameters. The resulting best fit and confidence interval band can also be observed in figure 2.
@@
The primary significance of this model is the accuracy in the predicted relationships between $Q$ and $f$ across three different regimes defined by competing dominant loss mechanisms. It is observed that thermoelastic damping limits device performance at cryogenic temperatures for some frequencies of interest, a characteristic that is not observed in quartz [18] due to the smaller diffusivity and phonon velocities when compared to the measured values for lithium niobate at low temperature [46]. Additionally, at high frequencies performance becomes limited by Rayleigh scattering of heavy impurity ions. This impurity loss represents a non-fundamental engineering barrier that could be overcome by better crystal purity, thus an additional idealistic loss model has also been plotted on figure 2 that shows the theoretical device performance of a pure crystal. Even assuming perfectly pure crystals, the intrinsic loss limit at 4 K in lithium niobate is still orders of magnitude below that of quartz resonators at the same temperature and in the presented frequency range. This is due to the stronger thermoelastic damping in lithium niobate when compared to quartz, which has been observed to follow the theoretical Landau-Rumer dissipation at this temperature and frequency range [35]. This observed limit for quartz is shown by the dashed black trace in figure 2 for direct comparison.
@@ VI Temperature Dependence of Acoustic Loss
The measurement of the quality factor’s temperature dependence, $Q(T)$ offers further insights into the dominant loss mechanisms, such as regimes where losses are primarily influenced by two-level systems (TLS) associated with impurity ions. By observing the scaling relationship of losses ($1/Q$) with temperature as shown for the B13,0,0 10.6 MHz mode on the main axes of figure 3, we see the presence of absorption peaks for defined temperature $\approx 28$ K and above. Such increased absorption in narrow temperature regions is due to the system entering an energy regime in which the thermal bath become resonant with phonons. These features have been observed in acoustic systems at low temperature previously [48], and are not indicative of TLS induced dissipation.
The characteristic scaling law expected for intrinsic losses due to the coupling of phonons to impurity TLS is theorised to follow $n=1/3$ [49, 18], however many experimental studies on low temperature acoustic systems [18, 50, 51, 52] have recorded a scaling slightly skewed from this exact value due to competing losses from multiple TLS ensembles and excess loss from other mechanisms such as resonant absorption. The regime between 10 and 20 K shows a temperature dependence of $1/Q\propto T^{0.34}$ that is reminiscent of TLS scattering. However, due to the presence of resonant absorption peaks in close proximity, and the loss saturation observed for $T<10$ K, it is hard to say with certainty if this effect pure TLS absorption, or multiple other competing mechanisms.
In the inset of figure 3, we show the further cooling of the $A_{19,0,0}$ 31 MHz mode down to $T=$10 mK, where a weaker temperature dependence was observed in the 1-4 K regime, most likely due to the presence of resonant absorption. At even lower temperatures the losses saturate and no further improvement is found, showing that the dominating dissipation channel cannot be ascribed to phonon-phonon scattering.
@@ VII Self Induced Transparency and Absorption
where $\omega$ is angular frequency, $\gamma$ parameterises linear dissipation, $F$ is the external driving force to the system, and $f(\dot{u}))$ is given by $f(\alpha)=-\beta_{1}\dot{\alpha}^{2}+\beta_{2}\dot{\alpha}^{4}$ with appropriate choices for the constants $\beta_{1,2}$.
This non linear model can then be solved utilising numerical harmonic balance methods in order to reproduce the anomalous self induced absorption effect. As the system is strongly-nonlinear, a variety of solution branches for $u$ are attained. In figure 5 a) the spectrum of normalised displacement amplitudes $u(f)$ is plotted for one of the solution branches, as a function of drive power $F$. The same symmetrical discontinuity about resonance can be observed as in figure 4, as the response collapses under self induced absorption due to the presence of non-linear dissipation that dominates above a threshold drive force. By visualising the parameter space for stable solutions to equation (4), plotted in figure 5 b), it is clear that the non-linear dissipation term causes switching between bi-stable states on both sides of the resonant peak. This recreates the characteristic self induced absorption phenomena. The accuracy of this model in reproducing the non-linear behaviour observed in figure 4 suggests that populations of heavy ions, to the level in parts per million, limit the performance of lithium niobate BAW resonators, as in un-swept quartz.

### v3

@@ I Introduction
Figure 1: a) Diagram of experimental set up. Cryostat environment was held under vacuum at temperatures of either 300 K or 4 K, and 10mK depending on the required measurement. The signal path includes attenuation $\alpha$ on the input line in order to reduce thermal noise on the drive signal, as well as an amplifier on the return line to maximise signal to noise ratio. b) Simulated mode distribution of the fifth order longitudinal mode. Red (blue) denotes the positive (negative) displacement in the $z$ direction
@@
Piezoelectric lithium niobate for example provides the vibrating media for tunable micro-ring resonators [22, 23], optical frequency combs [24], micro electromechanical system (MEMS) based resonators [25], devices for optical to microwave transduction [26, 27, 28, 29, 30] and acoustic quantum state operations [31, 32, 33]. This material presents an attractive candidate as a macroscopic BAW resonator due to a stronger piezoelectric coupling coefficient, and smaller loss tangent than crystalline quartz at room temperature. However, material properties display a strong temperature dependence, and are thus drastically different when cooled. This motivates further investigation into the performance of lithium niobate BAWs at cryogenic temperatures, for the purpose of next generation tests of fundamental physics. Additionally, the study of BAW modes at cryogenic temperatures allows one to investigate fundamental properties of the material.
In this work the loss mechanisms of lithium niobate BAW resonators are investigated by precise measurement of the quality factor ($Q$) of their acoustic modes at cryogenic temperatures. In the following sections II and III, the measurement set up is introduced and the frequencies of various fundamental bulk vibration modes are identified. Section IV then derives a theoretical model for the total $Q$ of a BAW resonator subject to accumulative loss mechanisms. In sections V and VI the measured $Q$ of various resonator modes are presented as functions of frequency and temperature respectively. Multiple competing loss mechanisms are identified based upon these observations. In section VII, we report on a strong drive amplitude dependence on the measured losses. Such self induced absorption and transparency effects are confirmed to be due to microscopic impurities by the use of an appropriate model of non-linear dissipation. Lastly, section VIII discusses the performance of these devices when further cooled to the extreme sub-Kelvin temperature of 10 mK.
@@ II Measurement Set Up
Figure 2: Transmission spectra of A and B polarisation modes are shown in figures a) and b) respectively, plotted for various modes in arbitrary logarithmic units normalised to the resonant peak. Insets in each figure show the same spectra in linear units with the corresponding fit that is used to extract linewidth $\gamma$.
@@
Figure 3: Measurement of a)$f\cdot Q$ product and b) intrinsic $Q$ as a function of frequency for various longitudinal modes in the crystal resonators at $T=4$ K. The measured intrinsic quality factors are represented in blue, the black trace and grey bands denote the best fit and 95% confidence interval to the model of equation (3). The individual influence of clamping, thermoelastic damping (TED) and Rayleigh scattering are shown by the dashed magneta, cyan and green traces respectively. The red trace shows the idealistic performance attainable without the presence of Rayleigh scattering. And the dashed black line gives the intrinsic Landau Rumer limit to losses observed in similar quartz devices [35].
In this configuration, as shown in figure 1, acoustic resonances can be excited by an electrical drive signal applied to the electrodes. By connecting each co-axial line to ports of a vector network analyser; the acoustic mode frequencies $f_{r}$ are identified as peaks of maximum power transmission $S_{21}$. Examples of such spectra are presented in figure 2. The input drive signal was varied across the frequency spectrum with 1 Hz resolution such that narrow resonances can be identified. The resulting spectra are then compared to a finite element model such that the mode structure of strongly coupled resonances can be identified.
@@ III Mode Structure
The frequencies of strongly coupled modes found by measurement were matched to the frequencies given by solutions to a finite element modelling as in figure 1 b), the results of which are presented in table 1. Close agreement between experimental data and model solutions is found. All odd overtones from $n=3$, to $n=15$, in both A and B polarisation are recorded in the table. For higher frequencies, the finite element resolution of the model needs to be further increased in order to guarantee accuracy. However from the results provided in table 1, a simple linear extrapolation $f_{i}/n_{i}=f_{j}/n_{j}$ can be utilised to identify any $n_{i}$ overtone, of frequency $f_{i}$, for modes of the same polarisation family given knowledge of another mode.
@@ IV Theoretical Model of Acoustic Loss
The dominant loss mechanisms in BAW resonators can be understood by measurement of the intrinsic quality factor of the acoustic modes $Q$. This quantity is the inverse of the total phonon losses and as such can be parameterised as the inverse sum of the quality factor associated with each competing loss mechanism. For bulk waves, the contributing losses to $Q$ can be due to clamping of the vibrational structure to external supports, thermoelastic damping, phonon-phonon scattering, and phonon-defect scattering. Phonon-electron scattering loss can be dominant in semi-conductor systems or at microwave frequencies, however it is negligible here.
For acoustic resonators at 300 K, phonon-phonon scattering loss is well described by the macroscopic Akheiser regime [36], which predicts a constant $Q\cdot f$ product. For lithium niobate the theoretical limit to $Q\cdot f$ as predicted by the Akheiser model is $Q\cdot f\approx 4.74\times 10^{14}$, although the largest room temperature value observed in the devices measured in this work is $1.12\times 10^{14}$ corresponding to a maximum $Q$ of $3.28\times 10^{6}$ However, at lower temperatures the life time of thermal phonons $\tau_{\mathrm{th}}$ increases such that phonon attenuation process undergoes a transition into the microscopic description of Landau and Rumer [37]. In this regime, the attenuation displays a weaker frequency dependence. As a result $Q\cdot f$ begins to scale as $Q\cdot f\propto f$, and $Q$ also greatly increases with decreasing temperature. It is precisely this change to the fundamental limit of intrinsic acoustic loss that gives rise to the numerous applications of high overtones BAW devices at low temperature. For lithium niobate at 4 K phonon modes are well described by the Landau-Rumer model for $f\gtrsim 1.4$ MHz, which equivalently gives $12\pi f\tau_{\mathrm{th}}>$ 1.
@@
where $\rho$ is the material density, $T$ is the reference temperature, $C^{s}$ is the specific heat capacity at constant strain, $c_{ij}$ is the isothermal elastic tensor and $\alpha_{ij}$ the thermal expansion tensor. The turning point frequency $f^{*}$ denotes the transition point between the adiabatic and isothermal propagation regimes. It is given by $f^{*}=V_{3}^{2}/2\pi\chi_{3}$ where $\chi_{3}$ is the material diffusivity. For lithium niobate at $T=4$ K the thermal conductivity and specific heat have been measured, but the elastic constants and expansion coefficients are unknown. Approximating the elastic constants from their room temperature values gives $f^{*}\approx 13$ MHz, with $f>f^{*}$ defining the isothermal regime.
@@
Acoustic attenuation due to either structural defects or microscopic impurities can be harder to quantify exactly, however scaling relationships can be predicted in some cases. Considering the standard tunnelling model of two level systems in amorphous glasses [40, 41], it can be shown [42] that the acoustic absorption by an ensemble of low lying two level systems states, potentially arising from non-crystalline impurities, results in a temperature dependent scaling of the acoustic losses $1/Q_{\mathrm{TLS}}\propto T^{1/3}$. At lower temperatures populations of these two level systems can become unsaturated and have a dominant impact on total acoustic dissipation.
In addition to low lying two level system energy states, Rayleigh-type scattering of phonons in metals and glasses off large impurity ions is predicted to cause attenuation that scales with acoustic wavelength $1/Q_{\mathrm{Rayleigh}}\propto f^{-3}$ [43]. This relationship however, has been observed for high overtones in similar quartz piezoelectric BAW systems [35], suggesting the presence of small populations of impurity ions that cause amorphous behaviour at high frequencies.
@@
$\frac{1}{Q}=\frac{1}{Q_{\mathrm{Clamping}}}+\frac{1}{Q_{\mathrm{TED}}}+\frac{1}{Q_{\mathrm{Rayleigh}}}+\frac{1}{Q_{\mathrm{TLS}}}+\frac{1}{Q_{\mathrm{Landau}}}.$ (3)
Where $Q_{\mathrm{Landau}}$ represents the intrinsic limit of phonon-phonon dissipation as per the Landau-Rumer model.
@@ V Frequency Dependence of Acoustic Loss
The intrinsic quality factor of bulk lithium niobate modes can be measured by fitting a Lorentzian model to the $S_{21}$ spectrum of a given mode and extracting its linewidth $\gamma$, such that $Q=f_{r}/\gamma$. The measurement uncertainty following this procedure is then quantified as the standard error associated to the fit parameters, such fits can be seen in the insets of figure 2. As the motional resistance of these modes are far greater than the source impedance of the electrical measurement network, the modes are severely under coupled, and thus the external measured quality factor $Q_{e}$ corresponds exactly to the intrinsic quality factor of the resonator $Q$. The measured values from this procedure are shown in figure 3. The cryogenic quality factors show consistent improvement from room temperature values across all frequencies, with a maximum recorded quality factor of $8.9\times 10^{6}$.
Assuming the first three terms of equation (3) dominate at $T=4$ K, and substituting equations (1), (2) and the functional form $Q_{\mathrm{Rayleigh}}=10^{a}f^{-3}$ , a fit can be performed to the measured quality factors as a function of frequency with $\lambda_{33}$, $\eta$ and $a$ as fit parameters. The resulting best fit and confidence interval band can also be observed in figure 3.
@@
The primary significance of this model is the accuracy in the predicted relationships between $Q$ and $f$ across three different regimes defined by competing dominant loss mechanisms. It is observed that thermoelastic damping limits device performance at cryogenic temperatures for some frequencies of interest, a characteristic that is not observed in quartz [18] due to the larger diffusivity and slower phonon velocities when compared to the measured values for lithium niobate at low temperature [46]. Additionally, at high frequencies performance becomes limited by Rayleigh scattering of heavy impurity ions. This impurity loss represents a non-fundamental engineering barrier that could be overcome by better crystal purity, thus an additional idealistic loss model has also been plotted on figure 3 that shows the theoretical device performance of a pure crystal. Even assuming perfectly pure crystals, the intrinsic loss limit at 4 K in lithium niobate is still orders of magnitude below that of quartz resonators at the same temperature and in the presented frequency range. This is due to the stronger thermoelastic damping in lithium niobate when compared to quartz, which has been observed to follow the theoretical Landau-Rumer dissipation at this temperature and frequency range [35]. This observed limit for quartz is shown by the dashed black trace in figure 3 for direct comparison.
@@ VI Temperature Dependence of Acoustic Loss
Figure 4: $Q$ as a function of temperature $T$ is measured for the B13,0,0 mode and the resulting scaling relationship between losses $1/Q$ and temperature is shown. Inset shows the scaling for the A19,0,0 mode when cooled to 10 mK. Temperature scaling relationships are identified, however of resonant absorption peaks dominate system behaviour.
The measurement of the quality factor’s temperature dependence, $Q(T)$ offers further insights into the dominant loss mechanisms, such as regimes where losses are primarily influenced by an ensemble of low energy two-level systems (TLS) associated with impurity ions. By observing the scaling relationship of losses ($1/Q$) with temperature as shown for the B13,0,0 10.6 MHz mode on the main axes of figure 4, we see the presence of absorption peaks for defined temperature $\approx 28$ K and above. Such increased absorption in narrow temperature regions is due to the system entering an energy regime in which the thermal bath become resonant with phonons. These effects are features of the Landau-Rumer model and have been observed in acoustic systems at low temperature previously. [48] They are not indicative of TLS induced dissipation.
The characteristic scaling law expected for intrinsic losses due to the coupling of phonons to impurity TLS is theorised to follow $n=1/3$ [49, 18], as has been observed in many experimental studies on low temperature acoustic systems [18, 50, 51, 52]. The regime between 10 and 20 K shows a temperature dependence $1/Q\propto T^{0.34}$ that has the correcting scaling to be due to TLS, however the presence of resonant absorption peaks in close proximity suggests that the most dominant loss mechanisms in the $T>10$ K regime are due to thermal phonon effects.
In the inset of figure 4, we show the further cooling of the $A_{19,0,0}$ 31 MHz mode down to $T=$10 mK, where a weaker temperature dependence was observed in the 1-4 K regime, most likely due to the presence of further resonant absorption. At even lower temperatures the losses saturate, showing that the dominating dissipation channel cannot be ascribed to phonon-phonon scattering. However, it is clear that these devices do not reach the regime of TLS loss limited behaviour as the low temperature saturation does not exhibit the appropriate scaling relationship ascribed to TLS dissipation.
@@ VII Self Induced Transparency and Absorption
where $\omega$ is angular frequency, $\gamma$ parameterises linear dissipation, $F$ is the external driving force to the system, and $f(\dot{u})$ is given by $f(\dot{\alpha})=-\beta_{1}\dot{\alpha}^{2}+\beta_{2}\dot{\alpha}^{4}$ with appropriate choices for the constants $\beta_{1,2}$.
This non linear model can then be solved utilising numerical harmonic balance methods in order to reproduce the anomalous self induced absorption effect. As the system is strongly-nonlinear, a variety of solution branches for $u$ are attained. In figure 6 a) the spectrum of normalised displacement amplitudes $u(f)$ is plotted for one of the solution branches, as a function of drive power $F$. The same symmetrical discontinuity about resonance can be observed as in figure 5, as the response collapses under self induced absorption due to the presence of non-linear dissipation that dominates above a threshold drive force. By visualising the parameter space for stable solutions to equation (4), plotted in figure 6 b), it is clear that the non-linear dissipation term causes switching between bi-stable states on both sides of the resonant peak. This bi-stability can facilitate the observed van-der-pol like oscillations in physical systems, and also recreates the characteristic self induced absorption phenomena. The general form of equation (4) has been shown previously to admit solutions in which the resonant peak recovers under self induced transparency [56]. Alternatively one can utilise a non-diverging non-linear dissipation function $f(\dot{u})$ in order to model self induced transparency.
The accuracy of this model in reproducing the non-linear behaviour observed in figure 5 suggests that small populations of heavy ions limit the performance of lithium niobate BAW resonators, as in un-swept quartz.